### v1

# ConCISE: Confidence-guided Compression in Step-by-step Efficient Reasoning

Abstract
Large Reasoning Models (LRMs) perform strongly in complex reasoning tasks via Chain-of-Thought (CoT) prompting, but often suffer from verbose outputs caused by redundant content, increasing computational overhead, and degrading user experience. Existing compression methods either operate post-hoc pruning, risking disruption to reasoning coherence, or rely on sampling-based selection, which fails to intervene effectively during generation. In this work, we introduce a confidence-guided perspective to explain the emergence of redundant reflection in LRMs, identifying two key patterns: Confidence Deficit, where the model reconsiders correct steps due to low internal confidence, and Termination Delay, where reasoning continues even after reaching a confident answer. Based on this analysis, we propose ConCISE (Confidence-guided Compression In Step-by-step Efficient Reasoning), a framework that simplifies reasoning chains by reinforcing the model’s confidence during inference, thus preventing the generation of redundant reflection steps. It integrates Confidence Injection to stabilize intermediate steps and Early Stopping to terminate reasoning when confidence is sufficient. Extensive experiments demonstrate that fine-tuning LRMs on ConCISE-generated data yields significantly shorter outputs, reducing length by up to ~50% under SimPO, while maintaining high task accuracy. ConCISE consistently outperforms existing baselines across multiple reasoning benchmarks.
ConCISE: Confidence-guided Compression in Step-by-step Efficient Reasoning
Ziqing Qiao1, Yongheng Deng1, Jiali Zeng2, Dong Wang1, Lai Wei1, Fandong Meng2, Jie Zhou2, Ju Ren1, Yaoxue Zhang1 1Tsinghua University 2Pattern Recognition Center, WeChat AI, Tencent Inc., China
1 Introduction
In recent years, Large Language Models (LLMs) have achieved remarkable progress in natural language understanding and generation, particularly in complex reasoning tasks. Large Reasoning Models (LRMs), such as OpenAI-o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025), which leverage the Chain-of-Thought (CoT) paradigm Wei et al. (2022), have demonstrated state-of-the-art performance on benchmarks involving mathematical problem solving and logical inference Muennighoff et al. (2025); Qwen Team (2024). However, a prominent limitation of LRMs is their tendency to generate excessively verbose reasoning chains Feng et al. (2025); Chen et al. (2024). This inefficiency impairs user experience and incurs considerable computational overhead, posing challenges for real-time applications and deployment in resource-constrained settings Team et al. (2025).
Figure 1: Our proposed method, ConCISE, achieves a more favorable trade-off between reasoning compression and task performance than existing methods.
To mitigate the verbosity of reasoning outputs, recent research has focused on compressing the reasoning chains generated by LRMs Qu et al. (2025); Sui et al. (2025). A mainstream and effective strategy is finetuning LRMs on concise reasoning datasets, enabling the LRMs to generate shorter reasoning chains directly during inference Ma et al. (2025); Chen et al. (2024); Munkhbat et al. (2025). The quality of such compression heavily relies on how these datasets are constructed. One common approach is sampling-based selection, which generates multiple candidate reasoning chains and selects the shortest correct one Team et al. (2025), or removes redundant reflection steps after the final answer is obtained Chen et al. (2024). While simple, these approaches lack control over generation and rely on the reasoning chains naturally generated by the model, which may still include unnecessary steps, reducing compression effectiveness. Another line of work relies on post-hoc pruning, where redundant or less informative reasoning steps are identified and removed from a complete reasoning chain Cui et al. (2025); Xia et al. (2025). However, such modifications can disrupt the continuity of reasoning, leading to degraded performance of LRMs after fine-tuning.
To address the limitations of existing approaches, we seek to construct reasoning chains that are both compact and logically intact, effectively removing redundant reflection steps without sacrificing the performance of LRMs after fine-tuning. To this end, we analyse the reasoning behavior of LRMs and identify that the reflection behavior is not solely determined by correctness, but is heavily influenced by the model’s internal confidence. In particular, insufficient confidence can cause the model to re-evaluate correct steps, resulting in redundant reflections. We thus adopt a confidence-guided perspective to explain why LRMs generate reflection steps. From this view, we identify two key patterns underlying redundant reflection: Confidence Deficit, where the model reconsiders correct intermediate steps due to low internal confidence; and Termination Delay, where reflection continues even after a confident final answer is reached. These patterns substantially inflate reasoning chains and offer actionable insights for constructing concise reasoning data.
Building on this analysis, we propose ConCISE (Confidence-guided Compression In Step-by-step Efficient reasoning)—the first framework that leverages a confidence-guided perspective to actively construct concise reasoning datasets by suppressing redundant reflection during generation. ConCISE consists of two complementary components—Confidence Injection and Early Stopping—designed to mitigate Confidence Deficit and Termination Delay, respectively. Confidence Injection inserts confidence phrases before potential reflection points to strengthen the model’s belief in intermediate steps and suppress unnecessary reflection. Early Stopping employs a lightweight confidence detector to monitor the model’s internal confidence and terminate reasoning once sufficient confidence is achieved. Through the synergy of these two mechanisms, ConCISE enables a more efficient and compact reasoning chain. We then fine-tune LRMs using ConCISE-generated data via SFT Wei et al. (2021) and SimPO Meng et al. (2024), and evaluate their performance on multiple reasoning benchmarks. Experimental results show that ConCISE achieves a superior trade-off between reasoning compression and task performance. As shown in Figure 1, ConCISE reduces average response length by ~50% under SimPO while maintaining high accuracy; under SFT, it also significantly outperforms existing methods in compression efficiency. These results demonstrate that LRMs fine-tuned with ConCISE-generated data learn to avoid generating redundant reflection steps, achieving efficient reasoning without sacrificing performance.
The main contributions of this work are summarized as follows:
  • We present the first confidence-guided explanation of redundant reflection in LRM reasoning, identifying key patterns that underlie reasoning verbosity.
  • We propose ConCISE, a framework that proactively suppresses redundant reflection during inference and constructs high-quality, concise reasoning data for LRM fine-tuning.
  • Through extensive evaluation on multiple reasoning benchmarks, we demonstrate that LRMs fine-tuned with ConCISE-generated data achieve a better trade-off between reasoning chain compression and task performance.
2 Related Work
Figure 2: The ConCISE reasoning pipeline applied to solving the equation $73x-2=$ 7. The figure compares the original 8-step reasoning process with the streamlined 5-step process generated by ConCISE. It illustrates how ConCISE models reflection based on confidence values, identifying Confidence Deficit and Termination Delay, and applying Confidence Injection and Early Stopping to suppress redundant reasoning. Here, $C_{i}$ denotes the internal confidence of each step, and $T_{i}$ represents its corresponding threshold.
Recent research has increasingly focused on mitigating verbosity and redundancy in reasoning chains generated by Large Reasoning Models (LRMs)Qu et al. (2025). While CoT enables step-by-step reasoning, it often produces long outputs that increase latency, computational costs, and may degrade final accuracyWu et al. (2025); Nayab et al. (2024); Wang et al. (2025). To address this, existing approaches can be broadly categorized into three types: input-based methods, decoding-time interventions, and model-based training strategies Sui et al. (2025).
Input-based Methods aim to promote concise reasoning by modifying input prompts without changing the model architecture or decoding strategyLee et al. (2025). Common techniques include imposing token limits or instructing the model to reason briefly Han et al. (2024); Renze and Guven (2024). However, these approaches often fail when the model does not consistently follow the instructions. Another approach, task routing, dynamically selects an appropriate model or reasoning strategy based on input characteristics like question complexity Ong et al. (2024); Chuang et al. (2025); Aytes et al. (2025). Its effectiveness, however, depends on the accurate assessment of input characteristics, which is not always reliable.
Decoding-based Methods intervene during inference to control the reasoning process. One direction compresses intermediate steps into latent representations rather than explicit text, which improves brevity but sacrifices interpretability Hao et al. (2024); Shen et al. (2025b). Another strategy uses dynamic decoding, where lightweight models or heuristics evaluate each reasoning step and decide whether to retain, modify, or discard it during generation Sun et al. (2024); Zhang et al. (2025); Xu et al. (2025). While effective at reducing verbosity, these methods introduce additional computational overhead and may disrupt the reasoning flow.
Model-based Methods train models to generate concise reasoning directly. One approach employs reinforcement learning with reward functions that penalize verbosity Shen et al. (2025a); Aggarwal and Welleck (2025); Luo et al. (2025); Arora and Zanette (2025), but such methods can be unstable, sensitive to reward formulation, and impose significant training costs. More commonly, LRMs are fine-tuned on datasets with concise reasoning chainsMa et al. (2025); Cui et al. (2025); Liu et al. (2024); Xia et al. (2025). These datasets are often created by selecting the shortest correct chains or removing redundant parts post-hoc. Although effective, these methods lack a principled understanding of redundancy, potentially retaining unnecessary steps or removing useful context, which may hurt compression efficiency or model performance.
Motivated by the limitations of existing methods, we propose a model-based framework ConCISE. Specifically, we achieve precise identification and suppression of redundant reflection steps throughout the entire reasoning process by analyzing and explaining reflection behavior from a confidence-guided perspective. This mechanism allows the model to generate concise and efficient reasoning chains while maintaining the flow of reasoning. By doing so, we construct a high-quality training dataset and effectively transfer this capability to the model through fine-tuning, enabling it to perform concise reasoning without compromising model performance.
3 Method
3.1 Confidence-guided Formulation of Reflection in LRM Reasoning
We observe that reflection in LRMs is not solely triggered by correctness judgments. In many cases, reflection steps are invoked even on correct reasoning steps, including those that have already been verified, leading to substantial redundancy. This suggests that reflection behavior is also linked to the model’s internal confidence rather than correctness alone. To explain this phenomenon, we adopt a confidence-guided perspective to formalize when and why LRMs engage in reflection.
Let $S_{i}={s_{1},s_{2},\ldots,s_{i}}$ denote the partial reasoning chain up to step $i$, where each $s_{j}$ is a textual reasoning unit. We associate each step $s_{i}$ with a confidence score $c_{i}\in[0,1]$, representing the model’s internal belief in the correctness of that step. The model’s generation policy $\pi_{\theta}$ maps the current reasoning context $S_{i}$ to the next step $s_{i+1}$.
To model the decision between proceeding and reflecting, we introduce a dynamic threshold $t_{i}\in[0,1]$, which may vary with the model architecture or the evolving context. At each step, the model compares its internal confidence $c_{i}$ with this threshold $t_{i}$ to determine whether to continue reasoning or enter reflection. Formally:
$s_{i+1}=\pi_{\theta}(S_{i})=\text{ReflectionStep}\quad\text{if }c_{i}<t_{i}$
Guided by this formulation, we analyze the reflective behavior of LRMs and formally characterize two recurrent patterns responsible for reflection-related redundancy: Confidence Deficit and Termination Delay. These patterns explain how the model’s internal confidence leads to the occurrence of redundant reflection. We will elaborate on each in the following subsection.
(a) Reflection probability after confidence injection.
(b) Detected confidence progression during reasoning.
Figure 3: Validation experiments for Confidence Injection and Termination Delay. The dataset used in these experiments is described in Appendix B.2.
3.2 Two Key Patterns of Redundancy
Confidence Deficit.
One major source of redundancy in LRMs stems from their tendency to undertrust their correct intermediate steps. LRMs often display unexpected hesitation despite exhibiting fine-grained reasoning capabilities and achieving higher stepwise accuracy, triggering reflection even on simple and unambiguous reasoning steps. This self-undermining behavior leads to redundant reflection with minimal semantic gain, ultimately increasing computational cost and disrupting the efficiency of the reasoning process. We refer to this phenomenon as Confidence Deficit. Formally, Confidence Deficit can be defined as the occurrence of a step $s_{i}$ satisfying:
$\exists i,\quad c_{i}<t_{i}\quad\text{despite }s_{i}\text{ being correct}.$
Termination Delay.
Beyond Confidence Deficit, LRMs exhibit another redundant reasoning pattern we term Termination Delay: After producing a confident final answer, the model is expected to conclude with minimal additional reasoning. However, it often continues to generate unnecessary reflection steps even after repeatedly verifying the same conclusion.
This behavior is attributed to a sharp increase in the confidence threshold $t_{i}$ after reaching the answer at step $i^{\prime}$, making it increasingly difficult for the model’s internal confidence $c_{i}$ to exceed $t_{i}$. As a result, even when $c_{i}$ becomes relatively high, it may still fall short of the heightened $t_{i}$, preventing the model from terminating, leading to verbose post-answer reasoning. Formally, Termination Delay can be characterized as:
$\exists i>i^{\prime},\;\forall j<i^{\prime},\quad t_{i}\gg t_{j}\;\text{and}\;c_{i}<t_{i}$
3.3 The ConCISE Framework
To mitigate the above patterns—Confidence Deficit and Termination Delay—we propose ConCISE, an integrated framework that dynamically steers the reasoning generation process to reduce redundant reflections and produce concise reasoning chains.
ConCISE integrates two core mechanisms: Confidence Injection, which proactively inserts confidence phrases to suppress unnecessary reflection steps during reasoning, and Early Stopping, which halts further generation once sufficient post-answer confidence is detected. Together, these mechanisms operate in synergy to generate concise reasoning chains without compromising inference quality. The overall process is illustrated in Figure 2.
3.3.1 Confidence Injection
To alleviate Confidence Deficit, we design a simple yet effective mechanism called Confidence Injection. The key idea is to boost the model’s internal confidence during reasoning by inserting designed phrases, referred to as confidence phrases, to prevent unnecessary reflection steps. Considering that indiscriminate injection could disrupt reasoning, we selectively insert confidence phrases at critical points where the model’s confidence is relatively low—that is, when $c_{i}<t_{i}$.
Specifically, at each reasoning step $s_{i}$, the model first generates the next step $s_{i+1}=\pi_{\theta}(S_{i})$, where $S_{i}$ denotes the reasoning chain up to step $i$. If $s_{i+1}$ is identified as a reflection step, indicating insufficient confidence, we retroactively modify the input by appending a confidence phrase $p_{i}$ sampled from a curated pool $\mathcal{P}$, and regenerate $s_{i+1}$. Formally, the updated generation process is defined as:
$s_{i+1}=\pi_{\theta}(S_{i})\quad\text{if}\;c_{i}\geq t_{i}\;\text{else}\quad\pi_{\theta}(S_{i},p_{i})$
We manually constructed the initial version of confidence phrase pool $\mathcal{P}$ based on observations of certainty in original reasoning chains. To ensure generalization, the pool includes diverse formulations, and confidence phrases are randomly sampled from it during inference. To evaluate and further refine the confidence phrase pool, we conducted experiments measuring the effectiveness of different phrases. Specifically, we inserted each candidate phrase at reflection positions and recorded the probability that the model still generated a reflection step. This probability indicates whether the updated confidence $c_{i}^{\prime}$ (after phrase insertion) exceeds the dynamic threshold $t_{i}$.
As shown in Figure 3(a), different phrases exhibit varying levels of effectiveness, some notably reduce the likelihood of reflection, while others have only a marginal impact. Notably, even the best-performing phrases still yield a reflection rate above 20%, suggesting that confidence injection reduces redundant reflections while retaining necessary verification. Based on these results, we selected the top 20 phrases with the lowest reflection rates to form the final confidence phrase pool. The composition of the phrase pool and further experimental details are provided in Appendix A.1.
3.3.2 Early Stopping
Although Confidence Injection effectively boosts the model’s confidence during reasoning, it does not fully mitigate the pattern of Termination Delay, which is caused by the significant increase in the dynamic threshold $t_{i}$ after the model generates an answer, leading the model to continue unnecessary reflections even when its confidence is already high. To address this, we design an Early Stopping mechanism based on direct estimation of the model’s internal confidence. Specifically, we construct a lightweight confidence detector to provide a quantitative proxy for the model’s internal confidence at each step after the first answer is generated.
We introduce a probing prompt (such as "So I’m") after the current reasoning context and analyze the probability distribution of generated continuations. Based on statistical analysis, we collect a set of confidence-indicative tokens $\mathcal{W}_{+}$, which reflect affirmations of high certainty. Formally, the detected confidence score $\hat{c}_{i}$ at step $i$ is calculated as:
$\hat{c}_{i}=\sum_{w\in\mathcal{W}_{+}}p(w\mid S_{i},\text{Probing prompt}),$
To validate the confidence detector and examine how model confidence evolves post-FAS(First Answer Step), we calculate the average detected confidence $\hat{c}_{i}$ across key reasoning stages, including FAS and the subsequent five reflection steps (Ref1 to Ref5). As shown in Figure 3(b), the model exhibits low confidence at the FAS, but after the first reflection, confidence increases sharply and continues to rise through subsequent reflections. Notably, the final reflection step has an average detected confidence of $0.93$, suggesting that for $i>i^{\prime}$, the dynamic threshold $t_{i}$ effectively stabilizes around this high value, motivating the introduction of a manually controlled lower threshold. Therefore, we define an early stopping threshold $t_{\text{e}}=0.5$. The model is instructed to terminate generation as soon as its detected confidence $\hat{c}_{i}$ exceeds $t_{\text{e}}$. Further details regarding the selection of $t_{e}$ and the confidence detector are provided in Appendix A.2.
In addition, we also compute the average and maximum $\hat{c}_{i}$ before the FAS (Avg Pre-FAS and Max Pre-FAS), both of which remain close to zero. This indicates that our confidence detector only becomes effective after the first answer is generated, ensuring that premature termination during early reasoning stages does not occur.
3.4 Building Efficient Reasoning Chains
ConCISE integrates Confidence Injection and Early Stopping to dynamically control the reasoning generation process. Given an input question, the LRM generates the reasoning chain step-by-step. At each step, the model first examines whether the next generated step is a reflection step. If a reflection is detected, Confidence Injection is applied by appending a confidence prompt and regenerating the step with reinforced internal confidence. After generating the next reasoning step, a lightweight confidence detector estimates the detected confidence score $\hat{c}_{i}$. If $\hat{c}_{i}$ exceeds a predefined threshold $t_{e}$, the model terminates generation, assuming sufficient confidence in the answer. Otherwise, the generated step is appended to the current reasoning chain, and the process repeats.. Algorithm 1 summarizes the overall coordinated reasoning process in ConCISE.
Algorithm 1 Algorithmic workflow of ConCISE for generating concise reasoning chains.
1:Preparation:
2:    Large Reasoning Model $\pi_{\theta}$
3:    Confidence Phrase Pool $\mathcal{P}$
4:    Early Stopping Threshold $t_{e}=0.5$
5:    Prompt Template $T$
6:Input: Question $Q$
7:Initialize: Reasoning chain $S_{0}=T(Q)$
8:for each reasoning step $i=1,2,\ldots$ do
9:     Generate $s_{i}=\pi_{\theta}(S_{i-1})$
10:     if $s_{i}$ is a reflection step then
11:         Apply Confidence Injection:
12:             Sample a confidence phrase $p_{i}\in\mathcal{P}$
13:             Regenerate $s_{i}=\pi_{\theta}(S_{i-1}+p_{i})$
14:     end if
15:     Compute detected confidence $\hat{c}_{i}$
16:     if $\hat{c}_{i}>t_{e}$ then
17:         Terminate
18:     end if
19:     Accept $s_{i}$ and update $S_{i}=S_{i-1}+s_{i}$
20:end for
3.5 Training Objective and Fine-tuning
To align LRM generation with the behavior encouraged by ConCISE, we first construct a fine-tuning dataset by applying the ConCISE pipeline to a set of reasoning questions. This yields concise reasoning traces in which redundant reflections are actively suppressed and timely termination is encouraged through the use of Confidence Injection and Early Stopping.
We fine-tune the model on this dataset using two standard learning paradigms: supervised fine-tuning (SFT), which directly imitates the revised reasoning chains, and Simple Preference Optimization (SimPO), which further amplifies the model’s preference for confident, efficient reasoning. Both approaches optimize a shared objective that governs the desired generation behavior:
$\begin{cases}\pi_{\theta}(S_{i})\rightarrow\pi_{\theta}(S_{i}+p_{i}),&\text{if}c_{i}<t_{i}\\ \pi_{\theta}(S_{i})\rightarrow\text{Terminate},&\text{if }i\geq i^{\prime}\text{ and }\hat{c}_{i}>t_{e}\end{cases}$
This objective encourages the model to learn a generation policy guided by internal confidence, rather than merely imitating surface text patterns. Specifically, when the model’s confidence $c_{i}$ at a given step falls below the threshold $t_{i}$, it learns to approximate the distribution conditioned on an injected confidence phrase $p_{i}$, which helps improve internal certainty and suppress unnecessary reflections. After the first answer is generated (at step $i^{\prime}$), if the detected confidence score $\hat{c}_{i}$ exceeds the early stopping threshold $t_{e}$, the model is explicitly trained to terminate reasoning at that point.
Through this process, the model learns to regulate its generation trajectory based on confidence, strengthening certainty during reasoning and halting once sufficient confidence is reached. As a result, the generation distribution favors precise and concise reasoning chains without compromising task performance.
4 Experiments
Model Method Math-500 GSM8K AIME24 GPQA Average
[…]
DS-7B Origin 90.8 3854.2 100% 93.1 1441.9 100% 54.2 13573.5 100% 51.0 8141.8 100% 72.3 100%
OverThinkSFT 92.2 2537.7 65.8% 93.0 1002.1 69.5% 52.5 11225.3 82.7% 51.1 7638.7 93.8% 72.2 78.0%
SpiritSFT 91.0 2934.6 76.1% 91.4 1107.4 76.8% 51.7 11528.7 84.9% 50.3 7084.3 87.0% 71.1 81.2%
ConCISESFT 92.0 2244.3 58.2% 92.9 831.9 57.7% 52.1 9750.8 71.8% 50.0 5891.5 72.4% 71.8 65.0%
OverThinkSimPO 91.4 2405.1 62.4% 92.9 879.4 61.0% 50.0 9603.2 70.8% 49.9 6304.9 77.4% 71.0 67.9%
SpiritSimPO 87.2 1765.1 45.8% 90.8 687.6 47.7% 38.3 6926.0 51.0% 50.2 5832.0 71.6% 66.6 54.0%
ConCISESimPO 91.0 1945.7 50.5% 92.1 715.1 49.6% 48.3 7744.7 57.1% 48.0 4859.4 59.6% 70.0 54.2%
DS-1.5B Origin 82.2 4783.7 100% 85.4 2219.0 100% 29.2 17464.9 100% 34.5 9492.3 100% 57.8 100%
OverThinkSFT 84.0 3295.6 68.9% 85.2 1199.7 54.1% 30.0 12893.3 73.8% 34.8 8600.0 90.6% 58.5 71.8%
SpiritSFT 83.8 3856.7 80.6% 84.8 1475.9 66.5% 28.8 13684.5 78.4% 32.3 8554.7 90.1% 57.4 78.9%
ConCISESFT 83.6 2700.8 56.5% 84.9 923.3 41.6% 30.0 11359.0 65.0% 35.0 7252.6 76.4% 58.4 59.9%
OverThinkSimPO 83.6 2737.8 57.2% 84.8 1002.9 45.2% 31.26 11465.1 65.7% 34.7 8249.8 86.9% 58.6 63.8%
SpiritSimPO 82.6 2455.3 51.3% 82.6 804.2 36.2% 26.66 9946.3 57.0% 33.1 6910.3 72.8% 56.2 54.3%
ConCISESimPO 83.6 2428.6 50.8% 84.3 803.0 36.2% 30.43 8810.1 50.4% 35.4 7056.1 74.3% 58.4 52.9%
Table 1: Evaluation results of OverThink, SPIRIT, and ConCISE fine-tuned via SFT or SimPO on DeepSeek-R1-Distill 7B and 1.5B, showing accuracy(Acc.), token count(Tok.), and compression ratio(CR) across four benchmarks.
Figure 4: Comparison of StepNum and StepLen across two models and four benchmarks. Stars indicate ConCISE-based methods, which achieve stronger stepwise compression than baselines while maintaining comparable accuracy.
4.1 Settings
Models. We evaluate ConCISE on two Large Reasoning Models (LRMs): DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-1.5B Guo et al. (2025). Although both models exhibit strong mathematical reasoning abilities, their substantial redundancy in reflection steps highlights the need for reasoning chain compression.
Question Set. We find that effective reasoning-chain compression can be achieved with relatively limited data. Based on this observation, we select 2,000 questions from the MATH training setHendrycks et al. (2021) for fine-tuning, ensuring that each question yields a correct answer under greedy decoding and that the set maintains diversity in difficulty levels and response lengths. Detailed analysis is provided in the AppendixB.1.
Evaluation Benchmarks. We evaluate model performance across four benchmarks: GSM8kCobbe et al. (2021), Math-500Hendrycks et al. (2021), AIME24Li et al. (2024), and GPQA_diamondRein et al. (2024). All experiments use the same decoding configuration with temperature = 0.6 and top_p = 0.95Guo et al. (2025). The maximum length is set to 16k for GSM8k, Math-500, GPQA_diamond, and 32k for AIME24. For AIME24 and GPQA_diamond, due to their higher difficulty and smaller sizes, we perform 8 sampling rounds per instance and report the mean values across all metrics to ensure stability and reliability.
Evaluation Metrics. We adopt three primary metrics to comprehensively assess both reasoning accuracy and compression effectiveness. Acc. denotes the final answer accuracy. Tok. refers to the average response length, measured in tokens. CR (Compression Rate) is defined as the ratio of the average response length to that of the original model, with lower values indicating higher compression.
Baseline Methods. We compare our method against existing approaches that aim to remove redundant reasoning and construct efficient reasoning datasets. OverThink.Chen et al. (2024) Samples eight responses from the LRM with a relatively high temperature and selects the shortest correct one. Only the first reflection step after the answer is retained. Spirit.Cui et al. (2025) Iteratively removes steps with minimal impact on perplexity (PPL), deleting 30% of the least critical steps. Adjacent contexts are merged when needed to maintain coherence.
Training Details. We construct training datasets based on the question set using ConCISE, OverThink, and Spirit. For Spirit, compression is applied to reasoning chains generated via greedy search. During training, we adopt two strategies: SFT and SimPO. SFT directly uses the question and its corresponding compressed reasoning trace for supervised fine-tuning. For SimPO, reject samples are also required Rafailov et al. (2023). Following the original OverThink setup, we use the longest correct reasoning chain among the eight samples as the reject samples. To ensure a fair comparison, ConCISE adopts the same strategy. For Spirit, the full reasoning chain before compression serves as reject samples. All models are trained with $lr=1e-6$; SFT runs for 2 epochs and SimPO for 1 epoch.
4.2 Reults and Analysis
This section evaluates multiple compression methods across various reasoning benchmarks. Our results show that ConCISE strikes a superior balance between compression and task performance compared to baseline methods, enabling models to effectively eliminate redundant reasoning steps. Additionally, models fine-tuned with ConCISE on mathematical datasets generalize well to GPQA_diamond, showcasing its robustness.
4.2.1 ConCISE achieves superior compression-performance balance
As shown in Table 1, ConCISE demonstrates strong reasoning chain compression under both SFT and SimPO settings while maintaining comparable accuracy. Under SFT, it achieves a compression rate of ~60%, and under SimPO, it further improves compression to ~50%.
In contrast, OverThink achieves competitive task performance but exhibits weaker compression under SFT. Although it achieves further compression under SimPO, its final compression rate and task performance are comparable to, or even slightly worse than, ConCISE under SFT. As for Spirit, it underperforms in both settings, achieving the weakest compression under SFT and suffering noticeable task performance degradation compared to the other two methods. While Spirit matches ConCISE’s compression rate under SimPO, its accuracy drops more severely, especially on challenging datasets like AIME24. Overall, ConCISE offers a better balance between compression and performance compared to baseline methods.
(a) Step-level analysis of reasoning chains on Math-500
(b) Comparison of training datasets of different methods
Figure 5: Extended analysis of reasoning chain characteristics. More details can be found in the Appendix B.3.
4.2.2 How does ConCISE achieve better compression-performance balance?
To deeply understand how ConCISE achieves a better balance between compression and performance, we perform a stepwise analysis of generated reasoning chains on benchmarks. As shown in Figure 4, we evaluate the average number of reasoning steps in the thinking stage(StepNum) and the average token count per step in the thinking stage(StepLen) across different benchmarks. ConCISE consistently achieves the fewest reasoning steps, while StepLen remains similar to that of the original model. Further details can be found in the AppendixA.3.
Further analysis, focusing on reflection steps in the generated reasoning chains on MATH-500, reveals how each method handles reflection behavior. We use several metrics, including StepNum, Non-RefNum, RefNum, Pre-FAS RefNum, and Post-FAS RefNum. Non-RefNum represents the number of non-reflection steps, and the last three metrics measure the number of reflection steps across the entire reasoning chain, before FAS(First Answer Step), and after FAS, respectively. The results are shown in Figure 5(a).
We find that, in terms of RefNum, Spirit-SFT retains the most reflection steps, followed by OverThink, while ConCISE and Spirit-SimPO retain fewer reflection steps. For Non-RefNum, OverThink and ConCISE perform similarly, while Spirit-SimPO slightly underperforms and also causes a performance drop. This suggests that none of the compression methods can effectively reduce non-reflection steps without compromising model performance. Consequently, it indicates that existing compression methods mainly achieve compression by pruning reflection steps.
For Pre-FAS RefNum, OverThink scores the highest, followed by Spirit, as neither method effectively controls early-stage reasoning. Spirit-SFT, on the other hand, has the highest Post-FAS RefNum, indicating its limited ability to identify and prune redundant steps. In contrast, ConCISE maintains consistent control over both pre- and post-answer reflections. These differences help explain the performance gaps: OverThink’s inability to prune pre-answer reflections limits its compression potential, while Spirit’s failure to identify redundant steps leads to suboptimal pruning, removing essential content and degrading performance. ConCISE, through confidence injection and early stopping, precisely removes redundant reflection steps without harming essential reasoning process.
4.2.3 How does the training dataset affect compression?
Since all methods share identical training configurations and hyperparameters, the quality of the training dataset directly determines the final compression performance. To further investigate which features of the training dataset are highly correlated with the final compression performance, we analyzed the reasoning datasets constructed by ConCISE, OverThink, Spirit, Greedy-Search (reject samples of Spirit), and Longest-of-N (reject samples of ConCISE and OverThink), with the evaluation metrics including Tok., StepNum, StepLen, RefNum, Pre-FAS RefNum, and Post-FAS RefNum. Figure 5(b) shows the distribution of these metrics across the training datasets.
Our results indicate that surface-level metrics such as Tok., StepNum, and StepLen do not directly correlate with the compression rate. For instance, ConCISE and OverThink have nearly identical values on these metrics, but ConCISE achieves much better compression after fine-tuning.
In contrast, other metrics reveal clear distinctions: OverThink records the highest Pre-FAS RefNum, while Spirit leads in Post-FAS RefNum. ConCISE, by comparison, consistently maintains lower values on both. This aligns with the reasoning chain analysis in the previous subsection, indicating that key features from the training data are effectively transferred to the fine-tuned model. These features enable the model to compress reasoning, particularly reflection steps, in line with the dataset design principles.
4.2.4 Can reasoning compression methods generalize across tasks?
Another important question is whether compression methods can generalize beyond the training data distribution. This property is crucial for the practical use of reasoning chain compression.Qu et al. (2025) In our setup, the training dataset is constructed from the question set of the MATH training set. Accordingly, MATH-500 serves as the in-distribution evaluation, GSM8K and AIME24 represent in-domain but out-of-distribution tasks, and GPQA_diamond serves as an out-of-domain task, differing substantially in content, format, and prompt structure.
As shown in Table 1, ConCISE consistently achieves a better balance between compression and performance than baseline methods across both in-domain and out-of-domain tasks. These results indicate that fine-tuning on efficient reasoning chains constructed by ConCISE enables models to structurally suppress redundant reflection steps, a property that holds regardless of task type or prompt format, without compromising overall performance.
Benchmark Method Acc. Tok. StepNum StepLen CR
Math-500 Origin 90.8 3854.2 113.7 33.9 100%
ConCISE 92.0 2244.3 63.0 35.6 58.2%
Conf. Inj. 91.8 2619.6 79.5 33.0 68.0%
Early Stop 92.4 2653.8 68.4 38.8 68.9%
GSM8K Origin 93.1 1441.9 41.4 34.8 100%
ConCISE 92.9 831.9 22.5 37.1 57.7%
Conf. Inj. 92.6 933.8 23.2 40.2 64.8%
Early Stop 92.7 1003.0 22.3 45.0 69.6%
AIME24 Origin 54.2 13573.5 463.0 29.3 100%
ConCISE 52.1 9750.8 333.0 29.3 71.8%
Conf. Inj. 51.3 10165.8 372.2 27.3 74.5%
Early Stop 54.2 12204.6 389.7 31.3 89.9%
Table 2: Ablation study of DeepSeek-R1-Distill-Qwen-7B across three mathematical benchmarks.
4.3 Ablation Study
We conduct an ablation study to evaluate the two core components of ConCISE: Confidence Injection and Early Stopping. For each mechanism, we construct a training dataset and fine-tune the base model using the same configurations as in the main experiments. All experiments are performed in the supervised fine-tuning (SFT) setting and evaluated across three mathematical reasoning tasks: Math-500, GSM8K, and AIME24. The results, summarized in Table 2, show that both mechanisms achieve similar reasoning accuracy to full ConCISE but exhibit weaker compression performance. Both methods focus on reducing redundant reflection steps while maintaining the integrity of the reasoning process.
Confidence Injection suppresses unnecessary iterations by reinforcing the model’s confidence, but it has a limited impact on terminating reasoning after the final answer. Early Stopping, on the other hand, halts excessive post-answer reasoning but does not address earlier stages. Interestingly, Early Stopping nearly preserves the original model’s accuracy on benchmarks like AIME24, suggesting that post-FAS reflection steps often contain substantial redundancy and can be safely stopped without harming task performance. In summary, Confidence Injection reduces redundancy across the entire reasoning sequence, while Early Stopping targets post-answer overthinking. Only by combining both can ConCISE achieve comprehensive redundancy elimination, resulting in superior compression without compromising reasoning quality.
5 Conclusion
In this work, we first modeled the generation of reflection steps in Large Reasoning Models (LRMs) from a confidence perspective. This modeling approach enabled us to systematically identify two primary patterns of redundancy: Confidence Deficit, where models unnecessarily undertrust intermediate steps, and Termination Delay, where reflection persists despite reaching a confident answer. To address these issues, we introduced ConCISE, a novel confidence-guided framework designed to enhance reasoning efficiency. ConCISE employs two targeted mechanisms: Confidence Injection to suppress redundant reflections by boosting internal confidence, and Early Stopping to halt generation upon detecting sufficient answer confidence. Our experiments empirically validate that ConCISE significantly compresses response length (~50% under SimPO) while preserving strong accuracy across diverse reasoning benchmarks. Notably, ConCISE achieves superior compression-accuracy trade-offs over existing methods and exhibits robust generalization to out-of-domain tasks. This study offers a principled approach to modeling and mitigating redundancy in LRM reasoning, enabling more resource-efficient and user-friendly AI systems.
Limitations
While ConCISE demonstrates strong performance in compressing reasoning chains, it is subject to certain limitations that merit further exploration:
Further Compression of Step Length.
While ConCISE effectively targets redundant reflection steps, analysis indicates substantial potential remains for intra-step compression based on average step length. Future work could explore integrating techniques like TokenSkip with ConCISE for greater compression.
Confidence Estimation During Reasoning.
Although ConCISE introduces a lightweight confidence detector for the post-answer phase, it currently lacks a direct mechanism to model confidence prior to the first answer, relying instead on reflection occurrences as an indirect proxy.
Model Dependency and Sensitivity.
Validated on DeepSeek models, ConCISE’s effectiveness and the optimal tuning of its components (e.g., confidence phrases, detector, thresholds) may be sensitive to model architecture or scale. Its performance across diverse LRMs requires further study, potentially necessitating model-specific adaptations.
Appendix A Details about ConCISE
A.1 Confidence Injection
The selection of appropriate confidence phrases is critical for the Confidence Injection component of ConCISE. We curated a pool of 20 distinct phrases, presented in Table 3, by observing the model’s native reasoning expressions, employing manual design, and performing experimental refinement (partly illustrated in Figure 3(a)). The purpose of these phrases is to inject confidence into the model, thereby preventing redundant reflection steps during reasoning. To mitigate potential overfitting to any specific phrase, one phrase is randomly selected from the pool for injection at runtime.
Table 3: The Pool of 20 Confidence Phrases used for Confidence Injection.
Index Confidence Phrase
1 Therefore
2 The reasoning holds
3 Previous steps are correct
4 All steps are valid
5 With this established
6 That sounds reasonable
7 Let’s go ahead
8 Alright, let’s carry on
9 Let’s proceed
10 Let’s progress
11 So, putting it all together
12 The logic stands firm
13 The reasoning process is valid
14 Good, let’s keep going
15 Everything seems reasonable so far
16 This part checks out
17 I think that’s solid. So
18 The reasoning holds, let’s keep going
19 Everything checks out, let’s move on
20 All steps are solid, let’s move forward
A.2 Early Stopping
probing prompt and confidence-indicated tokens. The effectiveness of the Early Stopping mechanism relies heavily on the accuracy of its confidence detector. To identify an appropriate probing prompt, we analyzed the reasoning chains generated by the DeepSeek-R1-Distill-Qwen-7B model using Greedy decoding on the training question set. This analysis revealed that the model frequently expresses confidence during the later part of the reasoning chain. Table 4 summarizes the frequencies of the most common self-reported confidence expressions observed within chains.
Table 4: Frequency of Common Confidence Expressions observed in 2000 Reasoning Chains.
Confidence Expression Frequency Count
I’m confident 423
I feel confident 170
I can be confident that 157
I can confidently say 70
This gives me confidence that 68
I’m pretty confident 43
I’m pretty sure 37
Based on the prevalence of phrases starting with "I’m" leading into confidence statements, we selected "So, I’m" as the fixed probing prompt. Subsequently, we identified "confident", "pretty confident", "sure", and "pretty sure" as the primary confidence-indicative tokens.
The confidence score, $\hat{c}_{i}$, for a given state $S_{i}$ is then calculated by summing the probabilities of generating these tokens or sequences immediately following the probing prompt "So, I’m":
[…]
Here, notation like $P(\text{token}\mid S_{i},\text{sequence})$ represents the conditional probability of generating the next token given the preceding state $S_{i}$ and the specific text sequence provided as context. As demonstrated in Figure 3(b) and supported by our experimental results, the confidence detector designed using this probing prompt and set of confidence-indicative tokens proves effective for the early stopping mechanism in ConCISE.
Early stopping threshold. The selection of the early stopping threshold, $t$, was guided by the principle of tailoring the reflection process to the apparent quality of the reasoning chain. Our objective was to configure the threshold such that:
  • Simple problems with rigorous reasoning steps require no post-answer reflection.
  • Problems of moderate difficulty benefit from exactly one round of reflection.
  • Difficult problems or those with less rigorous reasoning undergo two or more rounds of reflection.
To find a threshold value that aligns with this principle, we empirically evaluated $t_{e}\in\{0.4,0.5,0.6,0.7\}$. For each value, we measured the probability distribution of the stopping points across different stages: stopping after the initial answer generation (0 reflections), after the first reflection, after the second reflection, etc. The results of this analysis are illustrated in Figure 6.
Figure 6: Distribution of early stopping points (probability of stopping at Pre-FAS, FAS, Ref1, 2, …, 5) for different threshold values ($t_{e}$).
Observing the distributions in Figure 6, we found that setting the threshold $t_{e}=0.5$ yielded the stopping behavior that best matched our desired criteria across varying problem difficulties. Therefore, we selected $t_{e}=0.5$ as the definitive early stopping threshold for all subsequent experiments conducted in this work.
A.3 Reflection step detection
The detection of reflection steps, as mentioned throughout this paper, plays a crucial role in both the algorithmic flow and experimental evaluation of ConCISE. Specifically, we employ two distinct methods for identifying these steps, selecting the most suitable one depending on the specific application context. These methods and their corresponding usage scenarios are detailed below.
A.3.1 Rule-based detection.
Reflection steps generated by the model often exhibit distinct linguistic characteristics, frequently containing specific keywords (e.g., "wait", "alternatively") that signal potential reasoning shifts or reconsiderations. Consequently, a straightforward and effective strategy for identifying the start of such steps is rule-based detection. This method checks for the presence of predefined "reflection keywords" within the generated text of a given step, $S_{i}$.
[…]
where $k\subseteq s_{i}$ indicates that the keyword $k$ appears as a substring of step $s_{i}$.The set $K$ of reflection keywords used for rule-based detection consists of the following:
  • "wait", "alternatively", "check", "reconsider", "reflect", "rethink", "reconsidering", "reviewing", "reassess", "pause", "second thought", "reevaluate", "verify".
While effective for identifying the onset of a reflection sequence, this method typically cannot recognize subsequent steps within the same reflection process, as these later reflection steps often lack the initial triggering keywords. Despite this limitation, we utilize this rule-based approach within the Confidence Injection. Since Confidence Injection operates during the model’s generation process, identifying and potentially preventing only the first redundant reflection step is often sufficient to avoid generating the entire reflection. This method is particularly advantageous here due to its efficiency (introducing no computational overhead) and ease of implementation.
Similarly, this rule-based detection is used for the Figure 3(a) analysis to determine if an intervention (like inserting a confidence phrase) triggers the start of a new reflection, as only detecting the initial step with this lightweight method is necessary.
A.3.2 LLM-as-a-Judge
When the task requires identifying all reflection steps within an entire reflection behavior—a capability beyond the rule-based method—a more comprehensive approach is needed. For example, implementing the OverThink baseline necessitates isolating and retaining the complete first reflection sequence that occurs after the FAS(First Answer Step). To address this, we employ an LLM-as-a-Judge methodology.
Specifically, we utilize the Qwen-Max model as the judge. We provide it with both the original problem/question and the model’s complete reasoning chain as input. The prompt supplied to Qwen-Max includes precise definitions of the FAS and the characteristics defining a ’reflection behavior’. The model is instructed to return a structured output that annotates the input chain, identifying the index of the FAS and providing the indices for all detected reflection steps, grouped according to the reflection behavior they belong to. We then parse this structured output to finalize the detection and grouping of reflection steps. Illustrative examples of the prompt structure, input format, and expected response can be found in Figure 9, Figure 10, and Figure 11
Our observations indicate that Qwen-Max performs reliably in identifying FAS and grouping reflection steps when the reasoning chains are of moderate length. However, we noted a decline in performance for chains exceeding 5000 tokens. This degradation is likely attributable to the known limitations of current LLMs in processing very long contexts effectively. Consequently, for the implementation of the OverThink baseline comparison, reasoning chains longer than 3000 tokens were manually annotated to ensure accuracy.
Furthermore, the LLM-as-a-Judge approach is the basis for calculating several reflection-based metrics presented in our results, such as RefNum, Non-RefNum, pre-FAS RefNum, and post-FAS RefNum (shown in Figure 5(a) and Figure 5(b)).
Appendix B Details about Experiments
B.1 Question Set
Figure 7: Difficulty diversity of question set.
Figure 8: Length diversity of question set.
While the construction methodology for the question set is detailed in Section 4, here we focus on illustrating its diversity in terms of difficulty and length.Figures 7 and 8 demonstrate the diversity of the Question set. Difficulty distribution is derived from metric Level in the MATH dataset, while length distribution reflects reasoning chain lengths from DeepSeek-R1-Distill-Qwen-7B under greedy decoding.
B.2 Validation Dataset
For the experiments presented in Figure 3, we curated the Verification Dataset. This was necessary because the required analyses (including First Answer Step (FAS) and reflection detection, as shown in Figure 3(b)) rely on annotations from the LLM-as-a-Judge method (§A.3.2), whose reliability decreases on long reasoning chains. To ensure dependable results, the Verification Dataset comprises another 1000 reasoning chains selected from the MATH training set, filtered based on two criteria: yielding a correct answer under greedy decoding and having a reasoning chain length under 5000. This curated set facilitates reliable LLM-based annotation for the validation experiments.
B.3 Evaluation Dataset
The analysis presented in Figure 5 required LLM-as-a-Judge annotations. To ensure reliable results despite this method’s limitations with long contexts, we filtered the MATH500 dataset. We selected only those problems where reasoning chains generated by all twelve specified finetuning configurations (combinations of ConCISE/OverThink/Spirit methods, 1.5B/7B sizes, and SFT/SimPO techniques) were shorter than 5000 tokens. This yielded a common subset of 340 problems, on which our subsequent comparative analysis of the different methods was exclusively performed.
Prompt You are an AI assistant trained to analyze reasoning steps in a response. Your task has two parts: 1. Examine each reasoning step to determine if it’s part of a reflection process. 2. Identify the earliest step where the final answer (as later shown in boxed{}) is first derived, regardless of whether it is formally written or boxed at that moment. ### [Definition of Reflection]: 1. A reflection process is a sequence of one or more reasoning steps that recheck or doubt a previously made conclusion, such as double-checking calculations, using alternative methods. 2. Typical signals include (but are not limited to): ‘Wait’, ‘Alternatively’, ‘Just to double check’, ’But hold on’, etc. These signals usually mean the start of a new reflection process. 3. However, even without such phrases, if the content of a step reflects a verification or reevaluation, it should be marked as a part of a reflection process. ### [Output Format]: 1. Reflection Step: List all reflection processes as groups of steps. - If Step3 and Step4 form a reflection, write as (Step3, Step4) - If Step5, Step6, Step7 form a new reflection process together, list as a separate group: [(Step3, Step4), (Step5, Step6, Step7)] - Avoid putting a lot of steps into one single reflection process. 2. First Answer Step: Write the earliest step where the final answer is first derived(e.g., Step2). ### [Example]: Question: 2 + 3 = ? Response: Step1: I start with 2 + 3.  Step2: That gives me 5.  Step3: Wait, is that right? Let me make sure…  Step4: But hold on, let me double-check. Maybe I should…  Step5: Wait, no, I think…  Step6: Therefore, the result is 5  Step7: Alternatively, if I use…  Step8: So, the answer is 5  Step9: **Final Answer**: the answer is boxed5. Output: Reflection Step: [(Step3), (Step4), (Step5, Step6), (Step7, Step8)] First Correct Step: Step2 Explanations for Reflection Step: - There is ’wait’, a reflection keyword, in Step3, so Step3 is the start of a reflection process. - There is ’But hold on’ in step4, so step4 means the start of a new reflection process. - Then, there is ’wait’ in step5, so step5 means the start of a new reflection process. - There is ’therefore’ in step6, so (step5,step6) is in the same refleciton process. - There is ’alternatively’ in step7, so step7 is the start of a new reflection process. - There is ’so’ in step8, so (step7,step8) is in the same reflection process. ### [Task]: Now analyze the following question and response: f"Question: {question}" f"Response: {response}" Please output the reflection steps and the first correct step in the format without extra explanation.
Figure 9: Prompt of using the LLM-as-a-Judge method for reflection step detection.
Input Question: What day of the week will it be 40 days from Tuesday? Response: Step1: Alright, so I need to figure out what day of the week it will be 40 days from Tuesday. Hmm, okay. Let me think about how to approach this. Step2: First, I remember that each week has 7 days. So, if something happens every 7 days, it’s on the same day of the week next time. That might be useful here. Maybe I can figure out how many weeks are in 40 days and then see how many extra days that is. Step3: Let me do the division: 40 divided by 7. I think that’s 5 weeks and 5 days. Step4: Wait, let me check. 7 times 5 is 35, right? So, 40 minus 35 is 5. So, 40 days is equal to 5 weeks and 5 days. That makes sense. Step5: So, if today is a Tuesday, then 35 days later is also a Tuesday. Because 35 is exactly 5 weeks, which is 5 times 7. So, after 35 days, it’s still Tuesday. Now, I need to find out what day it is 5 days after Tuesday. Step6: Let me count: Tuesday plus one day is Wednesday, plus two days is Thursday, plus three days is Friday, plus four days is Saturday, and plus five days is Sunday. So, 5 days after Tuesday is $𝑆𝑢𝑛𝑑𝑎𝑦$. Step7: Therefore, 40 days from Tuesday is Sunday. Let me double-check that to make sure I didn’t make a mistake. If today is Tuesday and 7 days later is also Tuesday, then adding another 7 days each time cycles through the week. So, 35 days later, as we calculated, it is Tuesday again. Then, 5 days beyond that would be Sunday. Yeah, that seems right. Step8: Final Answer: The day of the week 40 days from Tuesday is $𝑆𝑢𝑛𝑑𝑎𝑦$.
Figure 10: An example input of using the LLM-as-a-Judge method for reflection step detection.
Output Output: Reflection Step: [(Step4), (Step7)] First Answer Step: Step6
Figure 11: An example output of using the LLM-as-a-Judge method for reflection step detection.